OpenStudy (anonymous):

Tell whether the system has one solution, infinitely solution, or no solution. 3x + y = -13 -2x + 5y= -54

OpenStudy (anonymous):

1. -3/1x - 13 2. 2/5x -270 Its solved just need to graph..

OpenStudy (anonymous):

One solution\[ \left(x=-\frac{11}{17},y=-\frac{188}{17}\right) \] Just solve the two equations using for example eliminations to find the result above.

OpenStudy (anonymous):

you want to put both equations in slope-intercept form. if both equations have the same slope but different intercepts, then there are no solutions. if both equations have the same slope and intercept (they are the exact same line), there are infinitely many solutions. If both equations have different slopes, the lines will intersect at one point only.